McGilla Golf has decided to sell a new line of golf clubs. The clubs will sell for $800 per set and have a variable cost of $400 per set. The company has spent $151,000 for a marketing study that determined the company will sell 78,000 sets per year for seven years. The marketing study also determined that the company will lose sales of 19,000 sets of its high-priced clubs. The high-priced clubs sell at $1,200 and have variable costs of $700. The company will also increase sales of its cheap clubs by 10,000 sets. The cheap clubs sell for $400 and have variable costs of $200 per set. The fixed costs each year will be $10,080,000. The company has also spent $1,411,000 on research and development for the new clubs. The plant and equipment required will cost $25,200,000 and will be depreciated on a straight-line basis. The new clubs will also require an increase in net working capital of $1,354,000 that will be returned at the end of the project. The tax rate is 36 percent, and the cost of capital is 14 percent.

The payback period is  years (Round your answer to 3 decimal places. (e.g., 32.161)), the NPV is $ (Negative amount should be indicated by a minus sign. Do not include the dollar sign ($). Round your answer to the nearest whole dollar amount. (e.g., 32)), and the IRR is  percent. (Do not include the percent sign (%). Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1
Cost of marketing study($151,000) and research & development ($1,411,000)are sunk cost and not relevant for this analysis
A Sales Price per set                          800
B Variable Costs per set                          400
C=A-B Contribution margin per set                          400
D Quantity of sales per year                     78,000
E=C*D Annual Contribution              31,200,000
F Contribution margin of high priced clubs 500 (1200-700)
G Quantity of lost sales of high priced 19000
H=F*G Loss of contribution per annum                9,500,000
I Contribution margin of Cheap clubs 200 (400-200)
J Quantity of Increase in sales of Cheap clubs                     10,000
K=J*I Addition to contribution per annum                2,000,000
L=E-H+K Total AnnualContributon              23,700,000
M Annual Fixed Costs              10,080,000
N=L-M Annual Profit (excluding depreciation)              13,620,000
P Cost of new equipment              25,200,000
Q=P/7 Depreciation per year                3,600,000
R=N-Q Before tax operating profit              10,020,000
S=R*(1-0.36) After tax operating profit(36% tax rate)                6,412,800
T=S+Q After tax annual operating cash flow              10,012,800
PV1 Present Value of annual operating cash flow $42,937,939 (Using PV function of excel with Rate=14%,Nper=7, Pmt=-T)
U Increase in net working capital in year 0 $1,354,000
PV2 Present Value of released working capital in year7                   541,109 (1354000/(1.14^7)
NPV=PV1+PV2-P-U NET PRESENT VALUE (NPV) $16,925,048
NET PRESENT VALUE $16,925,048
V=P+U InitialCash Flow                   (26,554,000)
T Annual Operating Cash Flow                     10,012,800
W=V/T PAY BACK PERIOD                                2.652 YEARS
YEAR 0 1 2 3 4 5 6 7
V Initial Cash Flow                   (26,554,000)
T Annual Operating Cash Flow                                    10,012,800    10,012,800    10,012,800    10,012,800    10,012,800    10,012,800    10,012,800
X Cash Flow due to release in working capital $1,354,000
Y=V+T+X NET CASH FLOW                   (26,554,000)                                    10,012,800    10,012,800    10,012,800    10,012,800    10,012,800    10,012,800    11,366,800
INTERNAL RATE OF RETURN (IRR) 32.75% (Using IRR function of excel on the net cash flow)